(1) Let;

  • A=the first row [-1 -1] second row [3 3]. Find all 2×2 materices, B such that AB=0

(2) Distinguish whether the given matrix is symmetric or not

  • A=the first row [2 1 3] second row [1 5 -3] third row [3 -3 7]
  • B=the first row[1 1 3] second row[1 2 2] third row[3 2 3]
1
Expert's answer
2020-09-14T19:18:05-0400

(1) Given "A=\begin{bmatrix}
 -1 & -1 \\
 3 & 3
\end{bmatrix}." Let "B=\begin{bmatrix}
 b_{11} & b_{12} \\
 b_{21} & b_{22}
\end{bmatrix}."

Then


"AB=\begin{bmatrix}
 -1 & -1 \\
 3 & 3
\end{bmatrix}\begin{bmatrix}
 b_{11} & b_{12} \\
 b_{21} & b_{22}
\end{bmatrix}="

"=\begin{bmatrix}
 -b_{11}-b_{21} & -b_{12}-b_{22} \\
 3b_{11}+3b_{21} & 3b_{12}+3b_{22}
\end{bmatrix}"

If "AB=0," then


"\begin{alignedat}{2}
 -b_{11}-b_{21}=0 \\
 -b_{12}-b_{22}=0 \\
3b_{11}+3b_{21}=0 \\
3b_{12}+3b_{22}=0
\end{alignedat}"

"\begin{alignedat}{2}
 b_{21}=-b_{11} \\
 b_{12}=-b_{22}
\end{alignedat}"

"B=\begin{bmatrix}
 c & -d \\
 -c & d
\end{bmatrix}, c,d\in \R"

(2)

"A=\begin{bmatrix}
 2 & 1 & 3 \\
 1 & 5 & -3 \\
 3 & -3 & 7
\end{bmatrix}=>A^T=\begin{bmatrix}
 2 & 1 & 3 \\
 1 & 5 & -3 \\
 3 & -3 & 7
\end{bmatrix}=A"

Then the matrix "A" is symmetric.


"B=\begin{bmatrix}
 1 & 1 & 3 \\
 1 & 2 & 2 \\
 3 & 2 & 3
\end{bmatrix}=>B^T=\begin{bmatrix}
 1 & 1 & 3 \\
 1 & 2 & 2\\
 3 & 2 & 3
\end{bmatrix}=B"

Then the matrix "B" is symmetric.